OpenStudy (anonymous):

How to solve this?? (2cos x + 1)(cos x + 1) = 0

OpenStudy (ajprincess):

(2cos x + 1)(cos x + 1) = 0(cos x + 1) = 0 means either 2cos x + 1= 0 or cos x + 1 = 0 can u solve nw?

OpenStudy (btaylor):

since they are 2 factors, the easiest way would be to set the factors equal to zero: 2 cos x + 1 = 0 --> cos x =-1/2 cos x + 1 = 0 -> cos x = -1 Then take the inverse cosines, and you get x = 120º (π/3) or x = 180º (π)

OpenStudy (anonymous):